Systems Staff Meeting Notes 12/15/03

Systems Staff Meeting Notes
December 15, 2003

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Horrible basketball, great volleyball games over the weekend.
  2. Water circle is expanding: now includes Marvin Crabb in addition to Martha and Carol Turner. Peter McKay is also inquiring.
  3. Bobbie’s brother’s dog ate some fertilizer pellets and may not make it. Dangerous stuff for pets.
  4. No TMT meeting last week due to the Christmas party.
Moves/Construction Status
  1. Room 1 is wired; staff members are in place. Some of the public machines are coming over today. Need to activate 3 additional switches and fiber pull.
  2. Completing the wiring in room 205 East (Special Collections). They want a third pull, but that will be difficult.
  3. MSL 107 – no confirmation from Facilities.
  4. Steve Shorb says that DOT will be ready to wire on 12/26/03. Bobbie checked with Judy – they are working at that time. Wiring crews will have access to work in the building. Bill asks if it will be set up for wireless? Are there barriers to circ functions with wireless? Fixed IP on wireless under wall plate may not work. Need to find out if Resource Services is already doing circ functions on their wireless laptops. Bill will try to track this info down.
  5. Projects still remaining: East 308, MSL 107 & Systems server room (AC, network wiring and power for Facilities security servers).
  6. Systems room 16 – Johnson MetaSys – will enable a jumper from the Mod Hub (ArcNet) to be the monitor. Facilities will use the room when they are not out at DOT.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Brad’s laptop has been ordered.
  2. Nothing major broken.
Networking
  1. Will – received calls this morning about Windows Update patch windows not responding – not going away. Reason this is occurring is unknown, but Will is working on it. It is patching some machines – but is failing in some machines as well.
  2. Trend patching is working fine unless machines are shut down, file/print sharing is turned off or the machine is not transferred to the correct server.
Software Projects Status
  1. SysRequest - Cynthia has the attachment feature completed and is working on the print function. Bill suggests rolling with the attachments and working on print later, if necessary.
  2. Storage Retrieval – Rich brought some changes – he discussed them with Cynthia and Hilton on Friday. They wanted 6 fields instead of 4 – Rich suggested 3 with text. Public interface – feeds into staff – Cynthia and Hilton will work on it tomorrow.
  3. Map GIS – Ying is working on it – lots of work with little improvement to show for it – they are limiting themselves and will run into problems if they want to upgrade because of these modifications.
  4. TILT & FRDI are done.
  5. DLC Tracking database – Ying working on it in C Sharp, but now Map GIS has a higher priority.
  6. CORE network disruption – public sign-ons should not have been disrupted. They need to see NT2R. Geoffrey set up EDUSER – getting localized. JUSER exists.
DP Coordinator
  1. Laura will be gone after Wednesday.
  2. Nicole, Nikia and a new student will start in January.


Back to the Systems Homepage

Last updated December 22, 2003
by Debra Fetzer
debharr@mail.uflib.ufl.edu